# Show Less, Instruct More: Enriching Prompts with Definitions and Guidelines for Zero-Shot NER

### v2

@@ 4.3 Compared Models
We compare SLIMER to several state-of-the-art approaches for zero-shot NER:
ChatGPT (Radford et al., 2019), prompted with the same strategy in Ye et al. (2023). It constitutes a baseline not specifically instructed for NER.
InstructUIE (Wang et al., 2023b), based on the flan-t5-xxl encoder-decoder model with 11B parameters.
UniNER (Zhou et al., 2023), a family of LLMs all based on the LLaMA-1-7B architecture. We evaluate the three variants: type, type+sup and def. The first is trained on full PileNER-type, described in Subsection 4.1. Additionally, type+sup. was trained also on a collection of human-labeled NER datasets. UniNER-def instead, distinguishes from type for the presence of short, automatically generated definitions of the entity type. These descriptive sentences differ from ours mainly because they are generated within the annotation process and therefore often vary for the same tag. They are also not as structured and detailed as those in SLIMER.
GoLLIE  (Sainz et al., 2024) is an LLM based on Code-LLaMA leveraging annotation guidelines formatted in a code-like representation. In the comparison we considered only the 7B version, since it has a number of parameters similar to most of the other approaches, SLIMER included.
@@
4.4 Results
Model Backbone #Params MIT CrossNER AVG
Movie Restaurant AI Literature Music Politics Science
ChatGPT gpt-3.5-turbo - 5.3 32.8 52.4 39.8 66.6 68.5 67.0 47.5
InstructUIE Flan-T5-xxl 11B 63.0 21.0 49.0 47.2 53.2 48.2 49.3 47.3
UniNER-type LLaMA-1 7B 42.4 31.7 53.5 59.4 65.0 60.8 61.1 53.4
UniNER-def LLaMA-1 7B 27.1 27.9 44.5 49.2 55.8 57.5 52.9 45.0
UniNER-type+sup. LLaMA-1 7B 61.2 35.2 62.9 64.9 70.6 66.9 70.8 61.8
GoLLIE Code-LLaMA 7B 63.0 43.4 59.1 62.7 67.8 57.2 55.5 58.4
GLiNER-L DeBERTa-v3 0.3B 57.2 42.9 57.2 64.4 69.6 72.6 62.6 60.9
GNER-T5 Flan-T5-xxl 11B 62.5 51.0 68.2 68.7 81.2 75.1 76.7 69.1
GNER-LLaMA LLaMA-1 7B 68.6 47.5 63.1 68.2 75.7 69.4 69.9 66.1
SLIMER w/o D&G LLaMA-2-chat 7B $46.4\pm{1.8}$ $36.3\pm{2.1}$ $49.6\pm{3.2}$ $58.4\pm{1.7}$ $56.8\pm{2.1}$ $57.9\pm{2.1}$ $53.8\pm{1.7}$ $51.3\pm{2.0}$
SLIMER LLaMA-2-chat 7B $\textbf{50.9}\pm\textbf{0.9}$ $\textbf{38.2}\pm\textbf{0.3}$ $\textbf{50.1}\pm\textbf{2.4}$ $\textbf{58.7}\pm\textbf{0.2}$ $\textbf{60.0}\pm\textbf{0.5}$ $\textbf{63.9}\pm\textbf{1.0}$ $\textbf{56.3}\pm\textbf{0.6}$ $\textbf{54.0}\pm\textbf{0.5}$
Table 3: Comparison of OOD performance for SLIMER and state-of-the-art models on MIT and CrossNER benchmark. With the exception of UniNER-def, all the competitors’ results are taken from their respective papers as listed in Section 4.3.
Model Backbone #Params $\mu$-Precision $\mu$-Recall $\mu$-F1 M-Precision M-Recall M-F1
UniNER-type LLaMA-1 7B 30.59 40.29 34.78 34.47 45.32 37.58
UniNER-def LLaMA-1 7B 25.00 51.29 33.62 24.66 48.22 31.80
UniNER-type+sup. LLaMA-1 7B 31.40 47.53 37.82 32.08 47.52 36.79
GoLLIE † Code-LLaMA 7B 28.82 26.63 27.68 27.53 22.56 24.13
GLiNER-L DeBERTa-v3 0.3B 42.55 19.31 26.57 41.16 22.16 24.34
GNER-T5 Flan-T5-xxl 11B 19.31 50.15 27.88 26.64 46.24 30.26
GNER-LLaMA LLaMA-1 7B 14.68 59.97 23.58 8.87 44.45 14.21
SLIMER w/o D&G LLaMA-2-chat 7B $44.00\pm{7.84}$ $37.52\pm{3.37}$ $40.41\pm{5.09}$ $39.74\pm{5.82}$ $38.49\pm{2.94}$ $35.90\pm{4.08}$
SLIMER LLaMA-2-chat 7B $\textbf{47.69}\pm\textbf{0.57}$ $\textbf{43.09}\pm\textbf{1.42}$ $\textbf{45.27}\pm\textbf{1.04}$ $\textbf{42.68}\pm\textbf{0.14}$ $\textbf{41.40}\pm\textbf{0.62}$ $\textbf{40.14}\pm\textbf{0.44}$
Table 4: Comparing SLIMER against state-of-the-art models on BUSTER to assess generalization capabilities on never-seen-before NEs. Models leveraging on guidelines are denoted with symbol †.
In order to reduce the overlap between training and test classes, we trained SLIMER on the PileNER-type subset described in Section 4.1, which has 391 distinct NEs. As already observed in literature, instruction tuning helps aligning the model to the tasks, and most of the gains in performance are achievable with little amounts of instructions (Wang et al., 2022a; Zhou et al., 2024; Zugarini et al., 2024). Hence, we only picked 5 examples per class from PileNER-type subset. In addition, for each example containing an annotated instance, we also included a negative example, i.e. one without any named entity. Overall, SLIMER was trained on 3910 input sequences, a small fraction of the data typically fed to other state-of-the-art models. To better understand the training set scales of different models, in Figure 5 we depicted them with circles proportional to their training sizes.
Out-of-distribution Domains.
Table 3 compares SLIMER against other state-of-the-art LLMs on MIT and CrossNER benchmarks. Differently from models like UniNER, GLiNER and GNER, the overlap between training and test named entity tags is significantly reduced (as shown in Table 2), thus operating in a more fair zero-shot scenario. Despite that, and the fact that we used only a fraction of the training data with respect to other models, SLIMER offers competitive performance, surpassing several of the existing state-of-the-art models. Moreover, our training data is entirely synthetic, whereas models like GoLLIE or UniNER-type+sup also exploit human-annotated labels. The importance of gold-annotated examples can be depicted by observing the 16% absolute increase between UniNER-type and UniNER-type+sup (see Table 3).
@@ 4.4 Results
To experiment the ability of existing models on never-seen-before labels, we extend the zero-shot evaluation on BUSTER, which is characterized by financial entities that are rather far from the more traditional tags observed by all models during training. For simplicity, we limited the evaluation to most performing approaches only, thus we omitted ChatGPT and InstructUIE in this experimentation. Results outlined in Table 4 exhibit an inverse trend with respect to OOD experiments. Indeed, best scoring state-of-the-art models in MIT and and CrossNER, such as the ones from GNER family, both under-perform in BUSTER, also due to their inability to work on long input texts  [^2] Except for GoLLIE, all the other approaches require input chunking into multiple smaller passages. While the sliding window for most is set to 900 words, for GNER we are limited to work on 150 words per input text.. Analogously, GLiNER and GoLLIE struggle in such a benchmark. Notably, GoLLIE’s performance is unexpectedly low, given that it is the only other to exploit annotation guidelines and it was fed with the same guidelines we provide to our SLIMER. SLIMER instead, appears to be the most effective in dealing with unseen labels, thanks to its lighter instruction-tuning methodology. Figure 5 clearly delineates such a behaviour.
Figure 5: Comparison of state-of-the-art models: F1 scores on unseen NEs in BUSTER (x-axis), OOD evaluation on MIT/CrossNER (y-axis). Circles’ size is proportional to the number of examples seen in training by each model.
@@
Table 5: Some examples of definition and guidelines. Absolute F1 gains between SLIMER and its version without definition and guidelines are reported. In red, we highlight performance degradation, in blue positive improvements between 0 and 10 and in green the F1 increases exceeding by over 10 points.
Increasing the number of training data.
We investigated how zero-shot NER performance change as the number of training instances increases. We proceeded in two directions: (1) increasing the number of unique NEs tags, while sampling a fixed total amount of samples per type; (2) varying the number of examples per NE, while keeping fixed the number of distinct categories. In particular, we set 10 examples (5 positive and 5 negative) per class in (1), and we set the number of distinct NEs to 50 in experiment (2). In both cases, we progressively add new classes/examples to the already selected ones. F1 scores are measured by considering MIT, CrossNER and BUSTER as a unique benchmark. By observing the results outlined in Figure 4, it emerges that increasing the number of unique entity types somewhat improves results, albeit most of the progresses are achievable with just 20 entity tags. Adding more than two examples per class brings little to no benefits, instead. This confirms what already observed in literature (Wang et al., 2022a; Zhou et al., 2024), further motivating our decision to instruct models on a small portion of the training data.
@@
As a final analysis, we assess the impact of definition and guidelines in SLIMER. We compare SLIMER with a version of it devoid of definition and guidelines, referred as SLIMER w/o D&G. Results, reported in tables 3,4 and Figure 4 consistently demonstrate how the definition and guidelines are helpful to the model. Indeed, there are improvements in F1 in both OOD and never-seen-before scenarios. Moreover, the absence of guidelines also significantly increase the standard deviation over multiple runs, thus demonstrating that definitions and guidelines also make the learning more consistent and stable. Notably, we can also observe from Figure 4 that in order to reach comparable performance, SLIMER w/o D&G requires more training data. To better understand how definition and guidelines contribute in improving the model, we show some qualitative examples in Table 5333Additional examples can be seen in Appendix A.. For each example, we report F1 scores obtained by using SLIMER or its version lacking of D&G. Information about an entity type not only helps in detecting novel NEs, like in the case of “Generic Consulting Company”, but it can also be beneficial to disambiguate polysemous tags, such as “trailer”. Moreover, specifications like the one in “politics”, can improve precision, avoiding tagging of instances not strictly belonging to the given class.
@@ 5 Conclusions
In this paper, we presented SLIMER, an instruction-tuned LLM for zero-shot NER. With a prompt enriched with definition and annotation guidelines, and a fine-tuning on a restricted set of entity tags, SLIMER, differently from most of the existing models, is specifically designed to better deal with unseen named entity tags. Experiments show that definition and guidelines steer the annotation process, especially on never-seen-before classes, thus yielding better predictions and a more stable learning. Furthermore, SLIMER performs comparably to state-of-the-art approaches, while being trained on a fraction of samples and entity types having little overlap with the test set.
In the future, we plan to broaden the scope of SLIMER to any Information Extraction problem. Moreover, we will investigate solutions to better scale instruction-tuned models for zero-shot NER to large set of entity classes to tag.
@@ Acknowledgments
Appendix A Comparing SLIMER to its baseline on some NE classes

### v3

@@ 4.3 Compared Models
We take into account the several state-of-the-art approaches for zero-shot NER:
ChatGPT (Brown et al., 2020), prompted with the same strategy in Ye et al. (2023). While it constitutes a baseline not specifically instructed for NER, it serves as a reference baseline to demonstrate the out-of-the-box performance of a pre-trained LLM, appropriately prompted for NER, as done in previous works Zhou et al. (2023); Zaratiana et al. (2023); Ding et al. (2024).
InstructUIE (Wang et al., 2023b), based on the flan-t5-xxl encoder-decoder architecture of 11B parameters.
UniNER (Zhou et al., 2023), a family of LLMs all based on the LLaMA-1-7B. Despite being an older backbone, authors themselves observed no significant improvements when experimenting their approach with LLaMA-2. Therefore, we keep their original model. We evaluate UniNER’s three variants: type is trained on full PileNER-type, described in Subsection 4.1; type+sup. was trained also on a collection of human-labeled NER datasets; def, on the other hand, differs from type by the presence of short definitions of the entity type, as described in Section 2.
GoLLIE  (Sainz et al., 2024) is a model based on Code-LLaMA leveraging annotation guidelines formatted in a code-like representation. In the comparison we considered only the 7B version, since it has a number of parameters similar to most of the other approaches, SLIMER included.
@@
4.4 Off-the-Shelf Models for Zero-shot NER
We compare SLIMER against state-of-the-art models on the benchmarks described in Section 4.1.
Out-of-Domain (OOD) inputs.
Table 3 compares SLIMER against state-of-the-art models on MIT and CrossNER benchmarks. We consider this benchmark as OOD, since most of the existing models have been trained on examples including most of the entity classes present in the test set (see last column and Table 2). However, our model has minimal training/test classes overlap, thus operating in a more zero-shot scenario. Despite that, and the fact that we used only a fraction of the training data with respect to other models, SLIMER offers competitive performance, surpassing several of the existing state-of-the-art models. Moreover, our training data is entirely synthetic, whereas models like GoLLIE or UniNER-type+sup also exploit human-annotated text. The importance of gold-annotated examples can be depicted by observing the 16% absolute increase between UniNER-type and UniNER-type+sup (see Table 3).
Model $\mu$-F1 M-F1
UniNER-type 34.78 37.58
UniNER-def 33.62 31.80
UniNER-type+sup. 37.82 36.79
GoLLIE † 27.68 24.13
GLiNER-L 26.57 24.34
GNER-T5 27.88 30.26
GNER-LLaMA 23.58 14.21
SLIMER w/o D&G $40.41\pm{5.09}$ $35.90\pm{4.08}$
SLIMER $\textbf{45.27}\pm\textbf{1.04}$ $\textbf{40.14}\pm\textbf{0.44}$
Table 5: Comparing SLIMER against SOTA models on BUSTER to assess generalization over never-seen-before NEs. Models leveraging on guidelines are denoted with symbol †.
@@ 4.4 Off-the-Shelf Models for Zero-shot NER
To experiment the ability of existing models on never-seen-before labels, we extend the evaluation on BUSTER, which is characterized by financial entities that are rather far from the more traditional tags observed by all models during training. For simplicity, we limited the evaluation to most performing approaches only, thus we omitted ChatGPT and InstructUIE in this experimentation. Results outlined in Table 5 exhibit an inverse trend with respect to OOD experiments. Indeed, best scoring state-of-the-art models in MIT and and CrossNER, such as the ones from GNER family, both under-perform in BUSTER, also due to their inability to work on long input texts [^3] Except for GoLLIE, all the other approaches require input chunking into multiple smaller passages. While the sliding window for most is set to 900 words, for GNER we are limited to work on 150 words per input text.. Analogously, GLiNER and GoLLIE struggle in such a benchmark. Notably, GoLLIE’s performance is unexpectedly low, given that it is the only other to exploit annotation guidelines and it was fed with the same guidelines we provide to our SLIMER. We speculate the CodeLLaMA backbone is less familiar with financial domain inputs and such long definition and guidelines. SLIMER instead, appears to be the most effective in dealing with unseen labels, thanks to its lighter instruction-tuning methodology. Figure 3 clearly delineates such a behaviour.
Figure 3: Comparing SOTA models: $\mu$-F1 scores on unseen NEs in BUSTER (x-axis), OOD evaluation on MIT/CrossNER (y-axis). Circles’ size is proportional to the number of examples seen in training by each model.
4.5 Comparing Zero-shot NER Approaches
Existing approaches are all trained under different conditions (e.g. backbone, training data, entities overlap), thus making difficult to evaluate the approach per se. To isolate the methodology from the rest, we compare them under the same conditions. We use the same backbone (LLaMA2-chat-7B) and fine-tune all the models as described in Section 4.2. Such a setting also allows a better evaluation of actual zero-shot performance of existing models. In this experiment, we selected the best two performing approaches, namely GNER and GoLLIE [^4] Further details are provided in Appendix B..
Results are reported in Table 4. We found that, without an extensive training corpus covering most of the entity classes, GNER performance drops of 18 F1 points, scoring now 6 points below SLIMER. Additionally, its fine-tuning required an increase in the number of samples per entity type from 5 to 50, further revealing its high dependency on training data. GoLLIE, on the other hand, better preserves its generalization capabilities to unseen entity types, scoring 1 point less than SLIMER in MIT and CrossNER. Moreover, the gap between SLIMER and GoLLIE in BUSTER further increases. Notably, GoLLIE’s performance decrease is not solely due to the new training corpus, but also to the use of a non-code-oriented backbone [^5] Using its original CodeLLaMA model would improve performance by 3%, as shown in Table 8, Appendix B.. This limits the applicability of such an approach to non-code-oriented LLMs, as noted by the authors themselves.
@@ 4.5 Comparing Zero-shot NER Approaches
Table 6: Some examples of synthetically generated definition and guidelines. Absolute F1 gains between SLIMER and its version without definition and guidelines are reported.
4.6 Ablation studies
We conducted ablation studies to understand how zero-shot NER performance is affected as training data size increases, with or without the inclusion of definition and guidelines.
Increasing training data.
We investigated how zero-shot NER performance changes as the number of training instances increases. We proceeded in two directions: (1) increasing the number of unique NEs tags, while sampling a fixed total amount of samples per type; (2) varying the number of examples per NE, while keeping fixed the number of distinct categories. In particular, we set 10 examples (5 positive and 5 negative) per class in (1), and we set the number of distinct NEs to 50 in experiment (2). In both cases, we progressively add new classes/examples to the already selected ones. F1 scores are measured by considering MIT, CrossNER and BUSTER as a unique benchmark. By observing the results outlined in Figure 4, it emerges that increasing the number of unique entity types somewhat improves results, albeit most of the progresses are achievable with just 20 entity tags. Notably, the poor scores achieved by the model when trained on very few examples, are not due to the model’s inability to identify entities, but rather to its struggles in aligning the output to the required format. Adding more than two examples per class brings little to no benefits, instead. This confirms what already observed in literature (Wang et al., 2022a; Zhou et al., 2024), further motivating our decision to instruct models on a small portion of the training data.
Figure 4: $\mu$ F1 scores of SLIMER and its baseline without D&G on MIT, CrossNER and BUSTER altogether, as we increase the number of unique NEs (top) and the number of samples per NE (bottom) seen in training.
@@ 4.6 Ablation studies
As a final analysis, we assess the impact of definition and guidelines. We compare SLIMER with a version of it devoid of definition and guidelines, referred as SLIMER w/o D&G. Results, reported in tables 3,5 and Figure 4 consistently demonstrate how the definition and guidelines are helpful to the model. Indeed, there are improvements in F1 in both OOD and never-seen-before scenarios. Moreover, the absence of guidelines also significantly increases the standard deviation over multiple runs, thus demonstrating that D&G also make the learning more consistent and stable. Notably, we can also observe from Figure 4 that in order to reach comparable performance, SLIMER w/o D&G requires more training data. To better understand how definition and guidelines contribute in improving the model, we show some qualitative examples in Table $6$Additional examples can be seen in Appendix A.. For each example, we report F1 scores obtained by using SLIMER or its version lacking of D&G. Information about an entity type not only helps in detecting novel NEs, but it can also be beneficial to disambiguate polysemous tags, such as “trailer”.
@@ 5 Conclusions
In this paper, we presented SLIMER, an instruction-tuned LLM for zero-shot NER. With a prompt enriched with definition and annotation guidelines, and a fine-tuning on a restricted set of entity tags, SLIMER, differently from most of the existing models, is specifically designed to better deal with unseen named entity tags. Experiments show that definition and guidelines steer the annotation process, especially on never-seen-before classes, thus yielding better predictions and a more stable learning. Furthermore, SLIMER performs comparably to state-of-the-art approaches, while being trained on a fraction of samples and entity types having little overlap with the test set. In the future, we plan to broaden the scope of SLIMER to any Information Extraction problem. Moreover, we will investigate solutions to better scale to large sets of entity classes to tag.
@@
Finally, in a real-world scenario, the three examples required to automatically generate definitions and guidelines might not be available for new unseen labels. Nevertheless, in practice, manually preparing three examples per entity class may be still easier than creating detailed definitions and guidelines for a human annotator.
@@ Acknowledgments
Appendix A Qualitative Analysis
@@
Figure 5: The 391 Named Entities in the PileNER-type subset, grouped by macro topics. “misc” (not shown) groups 26 NEs that do not fit into the defined topics.
Appendix B Further experiments
Model Backbone Guidelines MIT CrossNER BUSTER
Movie Restaurant AI Literature Music Politics Science AVG
GoLLIE CodeLLaMA-7B $\checkmark$ 53.9 45.4 54.9 56.8 66.5 64.2 58.0 57.1 14.5
GoLLIE LLaMA2-chat-7B $\checkmark$ 47.9 (-6.0) 39.2 (-6.2) 54.2 (-0.7) 51.4 (-5.4) 62.5 (-4.0) 61.9 (-2.3) 55.0 (-3.0) 53.2 (-3.9) 15.2 (+0.7)
Table 8: Instruction-tuning GoLLIE on the same sub-set of unique Named Entity types of SLIMER, using its original code-oriented LLM and a non-code-oriented LLM as the one employed by SLIMER.
GNER and GoLLIE training details.
To ensure reproducibility, we here detail the training setups used for GNER and GoLLIE, implemented using the same LLM backbone and subset of unique entity types of SLIMER. While GoLLIE fine-tuned effectively on same SLIMER’s training data, GNER required significantly more samples per entity (50 instead of 5) to achieve comparable performance. While adding more examples (100 per NE) did not improve performance, this highlighted GNER’s dependence on training data.
Re-implementing both GNER and GoLLIE approaches was easily done using their provided scripts and code hosted in their respective Github repositories. We kept the same training hyper-parameters in their original scripts, as described in their respective papers Ding et al. (2024); Sainz et al. (2024). We only increased the number of training epochs to 10, to align all the models with SLIMER.
Regarding GNER, because it’s not feasible to list all 391 NEs of PileNER-type subset in the prompt, for each input text we randomly sampled 10 entity types, ensuring that the positive tags in the passage of text were included. Regarding GoLLIE, we converted PileNER-subset to GoLLIE’s format by similarly sampling 10 total guidelines per input, ensuring the positive where included. We reduced the label noise to 0.2 and guidelines dropout to 0.1, because of the reduced number of training samples, while we did not rephrase the guidelines. Indeed, we provided the same definition and guidelines of SLIMER, formatting them as doc-strings.
GoLLIE-CodeLLaMA on PileNER-subset.
We conducted further experiments to the ones in Section 4.4 by implementing GoLLIE on the same training data as SLIMER, using its original code-oriented backbone. Results in Table 8 show a 4-point drop in performance when using a non-code-oriented backbone. This underlines a significant limitation in GoLLIE’s adaptability, as noted by Sainz et al. (2024) themselves, suggesting reduced applicability of the approach to LLMs that are not code-oriented.